OpenStudy (anonymous):

Find the solution of y = 6x + 1 for x = 5. A.(5,2/3) B. (–29, 5) C. (5, 36) D. (5, 31)

OpenStudy (hoblos):

y = 6x + 1 for x=5 y=6(5) +1 = 31 (5, 31)